OpenStudy (anonymous):

If g(x)+x sin(g(x)) =x^2 and g(1)=0, find g′(1). g′(1)= I keep getting 0 and I dont know where I'm going wrong. Please show step by step

OpenStudy (ash2326):

We have \[g(x)+x\sin (g(x))=x^2\] Let's differentiate this \[g'(x)+x\frac{d}{dx}(\sin g(x))+\sin g(x)\times \frac{d}{dx} x=\frac{d}{dx} (x^2)\] Simplifying now \[g'(x)+x\cos g(x)\times g'(x)+\sin g(x)=2x\] Do you get till here?

OpenStudy (anonymous):

why is it sin g(x) + sin g(x)? where did the second sin g(x) come from?

OpenStudy (anonymous):

ive got that far minus the extra sin g(x)

OpenStudy (ash2326):

It's because we have two functions x and sin (g(x)) \[\frac{d}{dx} p(x) q(x)\] I've applied the multiplicative rule of differentiation there \[\frac{d}{dx} p(x) q(x)=p'(x)q(x)+p(x)q'(x)\]

OpenStudy (anonymous):

oo ok so you dont have to use chain rule on the sin g(x) then. Is that because it's a single variable inside the sin? if it was sin g(2x) would you need to use the chain rule?

OpenStudy (ash2326):

We have to use chain rule, see the second term. Cos (g(x) * g'(x)

OpenStudy (anonymous):

right I hadnt looked at that part yet. So u did multiplication rule first then chain rule. Does it matter what order you do them in?

OpenStudy (ash2326):

Yes, first the differentiation of main function and then multiplication with the inner function's derivative

OpenStudy (anonymous):

ooo ok, thank you for your help!! :)

OpenStudy (ash2326):

No problem :)